OpenStudy (anonymous):

sinv=1/5. i need to find sin(2v)

jimthompson5910 (jim_thompson5910):

sin(v) = 1/5 sin^2(v) = (1/5)^2 sin^2(v) = 1/25 --------------------- sin^2(v) + cos^2(v) = 1 1/25 + cos^2(v) = 1 cos^2(v) = 1-1/25 cos^2(v) = 24/25 cos(v) = sqrt(24/25) cos(v) = sqrt(24)/sqrt(25) cos(v) = 2*sqrt(6)/5 -------------------------------------- sin(2v) = 2*sin(v)*cos(v) sin(2v) = 2*(1/5)*(2*sqrt(6)/5) sin(2v) = 4*sqrt(6)/25

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

heres a triangle if u need a picture